### v1

# Sharp systolic inequalities for invariant tight contact forms on principal $\mathbb{S}^{1}$-bundles over $\mathbb{S}^{2}$

Abstract
The systole $\text{sys}(\alpha)$ of a contact form $\alpha$ is defined as the shortest period of closed Reeb orbits of $\alpha$. Given a non-trivial $\mathbb{S}^{1}$-principal bundle over $\mathbb{S}^{2}$ with total space $M$, we prove that the inequality $\text{sys}(\alpha)^{2}\leq\text{Vol}(M,\alpha\wedge\mathrm{d}\alpha)$ holds true for any tight contact form $\alpha$ invariant under the corresponding circle action. Moreover, equality holds exactly for Zoll contact forms, namely, contact forms whose Reeb orbits are all closed and have the same minimal period. This implies a sharp systolic inequality for rotationally symmetric Finsler metrics on $\mathbb{S}^{2}$, and a particular case of a conjecture by Viterbo.
1 Introduction
A one-form $\alpha$ on a three-manifold $M$ is called a contact form when $\alpha\wedge\mathrm{d}\alpha$ is nowhere vanishing. It induces on $M$ the contact structure $\ker\alpha$, as well as a canonical vector field, namely the Reeb vector field $R$, defined by the following equations:
$\begin{cases}\iota_{R}\alpha=1\\ \iota_{R}\mathrm{d}\alpha=0.\end{cases}$
The flow $(\phi^{t})_{t\in\mathbb{R}}$ of $R$ is called the Reeb flow of $\alpha$. An important motivation for the study of Reeb flows comes from the fact that they generalize geodesic flows. Indeed, the cotangent bundle of any manifold can be equipped with a canonical one-form, called the Liouville form, whose restriction to the unit cosphere bundle induced by a Riemannian (or Finsler) metric on the base manifold is a contact form. The associated Reeb flow coincides with the cogeodesic flow of the metric. Moreover, the contact volume of this contact manifold is, up to a dimensional factor, the metric volume of the base manifold.
Remark 1.1.
If the context does not specify which contact form is being used, it will appear as an exponent on objects that depend on it.
1.1 The systolic ratio of a contact form
Given a compact Riemannian manifold $(M,g)$ of dimension $n$, we define the systole of $g$, which we write $\text{sys}(g)$, as the length of the shortest closed geodesic of $g$. On non-simply connected manifolds, one usually considers the homotopical systole, namely the length of the shortest non-contractible closed curve (necessarily a closed geodesic), but here we allow also contractible closed geodesics. The systolic ratio of $g$ is defined as $\rho(g)=\frac{\text{sys}(g)^{n}}{\text{Vol}(g)}$, where $\text{Vol}(g)$ is the Riemannian volume of $M$. The study of the systole and of the systolic ratio is the objective of systolic geometry (see for example [27] and [11] for an introduction to the subject). Among objects of interest in systolic geometry is the supremum of $\rho(g)$ when $g$ varies in the set of all Riemannian metrics on a given manifold. Computing the possible values of this supremum, understanding how they depend on the topology of the manifold and describing the metrics for which the supremum is attained are important and difficult questions.
Motivated by the Riemannian setting, the authors of [30] defined the systole of a contact form $\alpha$ on a manifold $M$, which we write $\text{sys}(\alpha)$, as the minimal period of a periodic orbit of the Reeb flow of $\alpha$. If $M$ is a closed orientable manifold, this number is finite since Taubes proved in [35] that Reeb flows on compact three-manifolds have a periodic orbit. The sytolic ratio of $\alpha$ is defined as the ratio $\rho(\alpha)=\frac{(\text{sys}\alpha)^{2}}{\text{Vol}(\alpha)}$, where $\text{Vol}(\alpha)$ is the contact volume of $M$. Note that if $\alpha$ is the contact form induced on a unit cotangent bundle by a Riemannian metric, the metric and contact notions of systolic ratio coincide up to a dimensional factor.
In general, the contact systolic ratio behaves quite differently than its Riemannian counterpart. For example, write $C(\xi)$ for the set of contact forms inducing a given co-orientable contact structure $\xi$ on $M$. It was proved that the contact systolic ratio is unbounded on $C(\xi)$, first when $M$ is the three-sphere and $\xi$ is the unique tight contact structure on $\mathbb{S}^{3}$ in [6], on compact three-manifolds in [7], and in higher dimensions in [34]. This has to be put in contrast with the fact that the Riemannian systolic ratio is known to be bounded on a large class of manifolds. Indeed, Gromov proved in [22] that the homotopical systolic ratio is bounded from above by a constant depending only on the dimension whenever the non-simply connected manifold is essential. The boundedness of the Riemannian systolic ratio for $\mathbb{S}^{2}$ was proved by Croke in [17]. There are no examples of compact manifolds for which the Riemannian systolic ratio is known to be unbounded. The study of the local maximizers of the systolic ratio on $C(\xi)$ was started in [30], and completed in [6] (on $\mathbb{S}^{3}$ for the standard contact structure), in [12] (in dimension 3) and in [2] (in any dimension), where it was proved that the local maxima of the systolic ratio are exactly Zoll contact forms, namely contact forms whose Reeb flow only has periodic orbits, all of them with same minimal period. Again, this has to be compared with the metric situation where maximizers of the systolic ratio are poorly understood.
1.2 Invariant tight contact forms and the main result
Notations
[…]
$\mathbb{S}^{1}$-principal bundles over a base $B$ are classified by homotopy classes of maps from $B$ to the classifying space of $\mathbb{S}^{1}$, which is $\mathbb{C}P^{\infty}$. This allows us to define the Euler class of an $\mathbb{S}^{1}$-principal bundle as the pullback of $-1\in H^{2}(\mathbb{C}P^{\infty},\mathbb{Z})$ (recall that $\mathbb{C}P^{\infty}$ is naturally oriented) by the associated map. Since $M$ is oriented, the principal bundle structure induces an orientation on the base $\mathbb{S}^{2}$ as follows: if $(X,Y;\Theta)$ is a positive local frame on $M$, then $(\pi_{*}X,\pi_{*}Y)$ is a local positive frame on $\mathbb{S}^{2}$. We then define the Euler number $e$ of the bundle as the integer given by pairing the fundamental class of $\mathbb{S}^{2}$ with the Euler class. The data of an orientation on $\mathbb{S}^{2}$ and of an integer $e$ determines a cohomology class in $H^{2}(\mathbb{S}^{2},\mathbb{Z})$, which by the classification of $\mathbb{S}^{1}$-principal bundles determines a unique $\mathbb{S}^{1}$-principal bundle over $\mathbb{S}^{2}$ with Euler number $e$ and canonically oriented total space. We give a geometric description of the Euler number in Remark 3.3.
In view of the fundamental dichotomy between tight and overtwisted contact structures in contact topology (see [19]), we focus on contact forms inducing a tight contact structure in this paper. We write $\Omega(e)$ for the set of $\mathbb{S}^{1}$-invariant positive tight contact forms on the total space $M$ of the above $\mathbb{S}^{1}$-principal bundle with Euler number $e$. In [21], Giroux proved the following results:
  • If $e<0$, then the contact forms in $\Omega(e)$ induce the unique (up to isotopy) positive contact structure that is transverse to the orbits of the action;
  • if $e\geq 0$, then the contact forms in $\Omega(e)$ induce a contact structure whose set of Legendrian $\mathbb{S}^{1}$-orbits is non-empty and connected.
We give below examples of tight $\mathbb{S}^{1}$-invariant contact forms on bundles over $\mathbb{S}^{2}$.
  • The standard contact form $\alpha_{0}=\frac{1}{4i}\sum\limits_{i=1}^{2}\left(\bar{z_{i}}\mathrm{d}z_{i}-z_{i}\mathrm{d}\bar{z_{i}}\right)$ on $\mathbb{S}^{3}\subset\mathbb{C}^{2}$ is invariant under the $\mathbb{S}^{1}$-action $t\cdot(z_{1},z_{2})\mapsto(e^{2\pi it}z_{1},e^{2\pi it}z_{2})$ given by its Reeb flow. More generally, any positive Zoll contact form on a manifold $M$ is invariant under the $\mathbb{S}^{1}$-action induced by its Reeb flow, and the induced principal bundle structure on $M$ has negative Euler number. Zoll contact forms on non-trivial $\mathbb{S}^{1}$-principal bundles over a symplectic manifold with integral symplectic form can be obtained by the Boothby–Wang construction, see Section 7.2 in [19].
  • Another class of examples come from a class of metrics on $\mathbb{S}^{2}$. A Finsler metric on a manifold $M$ is a continuous non-negative function $F\colon TM\rightarrow\mathbb{R}^{+}$ satisfying the following properties:
    • $F$ is smooth on the complement of the zero-section;
    • if $u\in TM$ is not zero then $\mathrm{d}^{2}F(u)(w,w)>0$ for all $w\in T^{\text{vert}}_{u}TM\setminus(\mathbb{R}u)$;
    • $F(\lambda v)=\lambda F(v)$ for $\lambda\geq 0$;
    • $F(v)>0$ if $v\neq 0$.
    Riemannian metrics are special cases of Finsler metrics, where $F$ is the Euclidean norm induced by the metric. As in the Riemannian setting, a Finsler metric induces a contact form $\alpha_{F}$ on its unit tangent bundle. We call a Finsler metric on $\mathbb{S}^{2}$ a metric of revolution if $F$ is invariant under a circle action on $\mathbb{S}^{2}$ that has exactly two fixed points and is free on their complement. This action lifts to a free $\mathbb{S}^{1}$-action on $T^{1}M$ the unit tangent bundle for $F$. The induced principle bundle structure induced on $T^{1}M$ has Euler number 2. The contact form $\alpha_{F}$ is invariant under that action. Hence, $\alpha_{F}$ is in $\Omega(2)$.
  • If $\alpha$ is a positive $\mathbb{S}^{1}$-invariant contact form on $M$ and $f\colon M\rightarrow]0,+\infty[$ is invariant under the same action, then $f\alpha$ is an $\mathbb{S}^{1}$-invariant positive contact form as well.
The main result of this work is a sharp systolic inequality for invariant tight contact forms on non-trivial $\mathbb{S}^{1}$-principal bundles over $\mathbb{S}^{2}$.
Theorem 1.2.
If $e\neq 0$, then for all $\alpha$ in $\Omega(e)$, we have
[…]
Moreover, equality holds if and only if $\alpha$ is Zoll.
Remark 1.3.
  • Since the set of invariant contact forms with fixed volume has infinite diameter for the Banach–Mazur pseudo-metrics on the set of contact forms introduced in [32] and [3], this result shows that boundedness of the systolic ratio is not a purely metric phenomenon, as the dichotomy “usually bounded metric systolic ratio”/“never bounded contact systolic ratio” might suggest.
  • We focus here on tight contact forms, for which we can prove the sharp systolic inequality above. However, the existence of a systolic inequality for invariant contact forms is not a purely tight phenomenon. We prove in a forthcoming paper a general systolic inequality for invariant contact forms on $\mathbb{S}^{1}$-principal bundles in dimension 3, valid for both tight and overtwisted invariant contact forms.
  • The statement of Theorem 1.2 leaves the case of invariant contact forms on $\mathbb{S}^{1}\times\mathbb{S}^{2}$ open. This will be an objective of further work.
1.3 Applications
1.3.1 Systolic inequality for Finsler metrics of revolution on $\mathbb{S}^{2}$
As discussed in [17], the supremum of the Riemannian systolic ratio for $\mathbb{S}^{2}$ is finite, larger than the one of the round sphere but its value is not known. The authors of [4] proved that the systolic ratio of a Riemannian sphere of revolution is maximized by Zoll metrics. They also extended this result to a special class of $\mathbb{S}^{1}$-symmetric Finsler metrics. We now explain how Theorem 1.2 allows us to extend this result to all $\mathbb{S}^{1}$-invariant Finsler metrics on $\mathbb{S}^{2}$.
A subtlety arising in Finsler geometry is that, as opposed to the Riemannian setting, there is no canonical notion of volume. As a matter of fact, several notions of volume coexist in the literature. As it is of symplectic nature, we will consider the Holmes–Thomson volume with the normalization convention of [4]: if $(S,F)$ is a Finsler surface, and $U$ an open subset of $S$, we define the Holmes–Thomson area of $U$ as
$\text{area}_{HT}(U)=\frac{1}{2\pi}\text{Vol}(T^{*1}U,\lambda_{\text{Liouville}}).$
In other words, $\text{area}_{HT}(U)$ is, up to a normalizing factor, the contact volume of the unit cotangent bundle over $U$ for the canonical Liouville form. We will write $\text{area}_{HT}(F)$ for the Holmes–Thomson area of a Finsler two-sphere $(\mathbb{S}^{2},F)$. More details on volumes in Finsler geometry can be found in [31]. We show that the following systolic inequality holds for any Finsler metric of revolution.
Corollary 1.4.
If $F$ is a Finsler metric of revolution on $\mathbb{S}^{2}$, then the inequality
[…]
holds, with equality if and only if $F$ is a Zoll metric.
Proof.
Let $F$ be a Finsler metric of revolution on $\mathbb{S}^{2}$. The contact form $\alpha_{F}$ is in $\Omega(2)$. Theorem 1.2 applied to $\alpha_{F}$ gives
$\text{sys}(\alpha_{F})^{2}\leq\frac{1}{2}\text{Vol}(\alpha_{F}).$
Since $\text{sys}(F)=\text{sys}(\alpha_{F})$ and $\text{Vol}(\alpha_{F})=2\pi\text{area}_{HT}(F)$, we get the desired inequality, with equality if and only if $F$ is Zoll. ∎
Remark 1.5.
The authors of [28] generalized the main result of [4] to a class of singular Riemannian metrics on $\mathbb{S}^{2}$ called rotationally symmetric spindle orbifold. In a similar fashion to Corollary 1.4, Theorem 1.2 extends the main result of [28] to rotationally symmetric Finsler metrics on spindle orbifolds, since the Reeb flow associated to such a metric lifts to a Reeb flow on $\mathbb{S}^{3}$ invariant under a free $\mathbb{S}^{1}$-action.
1.3.2 Viterbo conjecture for invariant domains
Another corollary of Theorem 1.2 is an application to a conjecture of Viterbo. Let $\omega_{0}=\sum\limits_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}$ be the standard symplectic form on $\mathbb{R}^{2n}$. We write Vol for the volume form induced by $\omega_{0}$. Viterbo stated the following conjecture in [36] (see for example [23] for an overview of the variations on this conjecture):
Conjecture 1.6 (Viterbo conjecture).
For any normalized symplectic capacity $c$ on $(\mathbb{R}^{2n},\omega_{0})$, and any bounded convex domain $U$ of $\mathbb{R}^{2n}$, we have
[…]
Moreover, equality holds if and only if $U$ is symplectomorphic to a ball.
Theorem 1.2 allows us to show that this conjecture holds in explicit situations, some of which are new.
Let $(\phi^{t})_{t\in\mathbb{S}^{1}}$ be a symplectic circle action on $(\mathbb{R}^{4},\omega_{0})$, and $U$ a smooth compact domain that is invariant under this action and with contact-type boundary. The action restricts to the boundary of $\partial U$, which is diffeomorphic to $\mathbb{S}^{3}$. If $Y$ is a Liouville vector field on an invariant neighborhood of $\partial U$, transverse to $\partial U$, then averaging $Y$ along the orbits of the action produces an invariant Liouville vector field $\tilde{Y}$, which is transverse to $\partial U$. The contraction $\iota_{\tilde{Y}}\omega_{0}$ then induces a tight contact form $\alpha_{U}$ on $\partial U$. Furthermore, $\alpha_{U}$ is invariant under the circle action restricted to $\partial U$. Whenever this action happens to be free on $\partial U$, Theorem 1.2 implies that
$\text{sys}(\alpha_{U})^{2}\leq\text{Vol}(\alpha_{U})=\text{Vol}(U),$
with equality if and only if $\alpha_{U}$ is Zoll, which by Proposition 4.3 in [6] is equivalent to the fact that $U$ is symplectomorphic to a ball.
For many normalized capacities $c$ it is known that $c(U)=\text{sys}(\alpha_{U})$ whenever $U$ is a smooth bounded convex domain. This is for example the case for the capacities $c_{HZ},c_{SH},c_{1}^{EH},c_{1}^{CH},c_{1}^{Alt}$, following the work of Ekeland, Hofer, Zehnder, Abbondandolo, Kang, Irie and Hutchings ([24] provides a related discussion). Following a consolidated tradition, we denote their common value on such domains by $c_{EHZ}$. As a consequence, Theorem 1.2 can be seen as a special case of Viterbo conjecture in dimension 4 for $c_{EHZ}$. This leads us to the corollary below.
Corollary 1.7.
Let $(\phi^{t})_{t\in\mathbb{S}^{1}}$ be a symplectic circle action on $(\mathbb{R}^{4},\omega_{0})$, and $U$ a smooth convex domain that is invariant under this action. If the restricted action on $\partial U$ is free, then we have
[…]
with equality if and only if $U$ is symplectomorphic to a ball.
As a concrete example, let $\alpha_{0}=\frac{1}{2}\sum\limits_{i=1}^{2}x_{i}\mathrm{d}y_{i}-y_{i}\mathrm{d}x_{i}$ be the radial Liouville form on $\mathbb{R}^{4}$. Let $(\Phi_{t}^{+})_{t\in\mathbb{S}^{1}}$ and $(\Phi_{t}^{-})_{t\in\mathbb{S}^{1}}$ be the two circle actions on $\mathbb{C}^{2}$ defined by
$\Phi^{+}_{t}(z_{1},z_{2})=(e^{2i\pi t}z_{1},e^{2i\pi t}z_{2})\hskip 28.45274pt\text{ and}\hskip 28.45274pt\Phi^{-}_{t}(z_{1},z_{2})=(e^{2i\pi t}z_{1},e^{-2i\pi t}z_{2}).$
Identifying each $(x_{i},y_{i})$-plane of $\mathbb{R}^{4}$ with a copy of $\mathbb{C}$ yields two symplectic actions on $\mathbb{R}^{4}$. Moreover, those actions are free on $\mathbb{R}^{4}\setminus\{0\}$. Hence, as implied by Corollary 1.7, any convex domain $U$ that is invariant under any of those action satisfies the Viterbo conjecture. We have respectively $e=-1$ and $e=1$ for the induced actions of $\Phi^{+}$ and $\Phi^{-}$ on $\partial U$. As we will see from Proposition 2.4, this makes Corollary 1.7 essentially trivial for $\Phi^{+}$, but non-trivial and new for $\Phi^{-}$. The non-triviality of the result for $\Phi^{-}$ can be illustrated by the fact that the restriction of $\Phi^{-}$ to the unit sphere in $\mathbb{C}^{2}$ is a lift to $\mathbb{S}^{3}$ of the circle action on $T_{1}\mathbb{S}^{2}=\mathbb{R}P^{3}$ described in 1.3.1, which follows from the classification of $\mathbb{S}^{1}$-principal bundles over $\mathbb{S}^{2}$ by the Euler number. As such, Corollary 1.4 can be seen as a consequence of Corollary 1.7 for the action $\Phi^{-}$.
1.4 A sketch of the proof
The proof of Theorem 1.2 splits into two cases, depending on the sign of the Euler number of the bundle. When it is negative, the result is proved by an easy argument, see for example [25]. This is the content of Proposition 2.4.
When the bundle has positive Euler number and the invariant contact structure has a non-empty connected set of Legendrian $\mathbb{S}^{1}$-orbits, the proof goes as follows.
  • The first technical tool coming into play is a well-suited choice of surface of section for the Reeb flow of an invariant contact form. Given a flow on a manifold $M$, a surface of section for this flow is an embedded surface $S\subset M$ transverse to the flow, whose boundary is a union of orbits of the flow, and with the property that for any $x$ in $S$, the orbit of $x$ by the flow hits $S$ again in forward time. Knowing that such a surface exists allows us to reduce the study of the flow, or at least part of the flow, to the study of the first-return map defined on $S$, and hence to reduce the dimension of the problem by one. The use of surfaces of section in dynamics dates back to Poincaré in his study of the three-body problem, and has been very fruitful since then (see for example the surveys [20] and [33]). It has proved to be helpful in the quantitative study of Reeb flows (see for example [9], [29], [13], [8], [1], [16], [14], [15], [18] for recent applications) and in particular for proving sharp systolic inequalities in both the metric and contact setting (see for example [5], [6], [34], [4], [12], [10]). We construct in Section 3.1 an invariant surface of section $\Sigma$ for the Reeb flow of any $\mathbb{S}^{1}$-invariant contact form. This surface of section is a union of embedded cylinders, each of them invariant under the $\mathbb{S}^{1}$-action. Even though $\Sigma$ has a priori several connected components, we will only make use of the one that intersects the set of Legendrian $\mathbb{S}^{1}$-orbits.
  • By $\mathbb{S}^{1}$-invariance of the surface of section and of the Reeb flow, studying the first-return map boils down to the study of a real-valued function, which we obtain using the decomposition of invariant contact forms obtained in Section 3.2. This function both encodes information on periodic orbits of the Reeb flow and their periods, and on the contact volume of $M$ via its integral. Building on techniques used in [4], we use this function to find a suitable $2|e|$-periodic point of the first-return map, corresponding to a periodic Reeb orbit with period less than $\sqrt{\frac{\text{Vol}(\alpha)}{|e|}}$. This yields the result.
1.5 Acknowledgments
The author is very grateful to Alberto Abbondandolo and Rémi Leclercq for numerous stimulating conversations and advice. The author is partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 281071066 – TRR 191 and by the ANR grant 21-CE40-0002 (CoSy).
2 Generalities on $\mathbb{S}^{1}$-invariant contact forms
We introduce in this section the first integral of the Reeb flow induced by the $\mathbb{S}^{1}$-symmetry, and give some of its properties. It is a general fact that to a symmetry of a dynamical system corresponds an invariant of the dynamics. In our case, this invariant is given by the moment map of the action.
Definition 2.1.
Let $\alpha\in\Omega(e)$. The moment map of $\alpha$ is the function $\tilde{K}\colon M\rightarrow\mathbb{R}$ defined by
$\tilde{K}=i_{\Theta}\alpha.$
By $\mathbb{S}^{1}$-invariance, $\tilde{K}$ induces a function $K$ on $\mathbb{S}^{2}$. The following proposition states the standard fact that $\tilde{K}$ is indeed an invariant of the Reeb flow.
Proposition 2.2.
$\tilde{K}$ is left invariant by the Reeb flow.
Proof.
We must prove that $\mathcal{L}_{R}\tilde{K}=0$, which follows from Cartan’s formula:
$\mathcal{L}_{R}\tilde{K}=(i_{R}\mathrm{d}+\mathrm{d}i_{R})i_{\Theta}\alpha=i_{R}\mathrm{d}i_{\Theta}\alpha=i_{R}\mathcal{L}_{\Theta}\alpha-i_{R}i_{\Theta}\mathrm{d}\alpha.$
Since $\alpha$ is $\mathbb{S}^{1}$-invariant, one has $\mathcal{L}_{\Theta}\alpha=0$, hence the first term is zero. The second term vanishes as $R$ is in the kernel of $\mathrm{d}\alpha$. ∎
An important property of the moment map is that there is a bijection between the critical points of $K$ and the Reeb orbits of $\alpha$ that coincide with an $\mathbb{S}^{1}$-orbit as sets. Moreover, the period of those Reeb orbits is up to a sign the corresponding critical value of $K$.
Proposition 2.3.
The differential of $\tilde{K}$ is $\mathrm{d}\tilde{K}=-i_{\Theta}\mathrm{d}\alpha$. Moreover, the critical points of $K$ are in bijection with $\mathbb{S}^{1}$-orbits of critical points of $\tilde{K}$, and themselves are exactly the $\mathbb{S}^{1}$-orbits that coincide as sets with Reeb orbits. The associated critical value of $K$ is the minimal period of this Reeb orbit up to sign, depending on whether the orientations of the corresponding Reeb orbit and $\mathbb{S}^{1}$-orbit coincide or not.
Proof.
We can apply Cartan’s formula to compute $\mathrm{d}\tilde{K}$:
$\mathrm{d}\tilde{K}=\mathrm{d}i_{\Theta}\alpha=\mathcal{L}_{\Theta}\alpha-i_{\Theta}\mathrm{d}\alpha=-i_{\Theta}\mathrm{d}\alpha$
which proves the first claim. We obtain from $\mathrm{d}K\circ\mathrm{d}\pi=\mathrm{d}\tilde{K}$ that $\mathrm{d}K$ vanishes at $x\in\mathbb{S}^{2}$ if and only if $\Theta_{y}$ is in the kernel of $\mathrm{d}\alpha_{y}$, with $\pi(y)=x$. Since the kernel of $\mathrm{d}\alpha$ is spanned by $R$, $\mathrm{d}K$ vanishes at $x$ if and only if $\Theta_{y}$ and $R_{y}$ are parallel. By $\mathbb{S}^{1}$-invariance of $R$ and $\Theta$, this happens if and only if $\mathbb{S}^{1}\cdot y$ and the Reeb orbit starting at $y$ coincide as set. In that case, writing $\gamma$ for the Reeb orbit starting at $y$, with minimal period $T$, and $\gamma_{\Theta}$ for $\gamma$ parameterized by the flow of $\Theta$, one has
$T=\int_{\gamma}\alpha=\left|\int_{\gamma_{\Theta}}\alpha\right|=\left|\int_{0}^{1}\alpha_{t.y}(\Theta)\mathrm{d}t\right|=|\tilde{K}(\gamma)|=|K(x)|.$
A corollary of Proposition 2.3 is that 00 is always a regular value of $K$. Indeed, $K$ vanishes exactly when $\Theta$ lies in the contact structure, hence $\Theta$ and $R$ cannot be parallel when $K$ vanishes. We finish this section by proving the inequality stated in Theorem 1.2 when the Euler number is negative. For a tight invariant contact structure, this happens exactly when the action of $\mathbb{S}^{1}$ is transverse to the contact structure. In that specific case, this inequality was already proved in the blog post [25] of Hutchings (see also Proposition 1.4 in [23]).
Proposition 2.4.
Let $e<0$ and $\alpha\in\Omega(e)$. Then the inequality
[…]
holds. Moreover, equality holds if and only if $\alpha$ is Zoll.
Proof.
Since $e<0$, we know that the $\mathbb{S}^{1}$-action is transverse to the contact structure. As a consequence, $\tilde{K}^{\alpha}$ does not vanish. This means that $\alpha_{0}=\frac{\alpha}{\tilde{K}^{\alpha}}$ is an $\mathbb{S}^{1}$-invariant contact form on $M$, and $\tilde{K}^{\alpha_{0}}=1$. Since $\tilde{K}^{\alpha_{0}}$ is constant, all the Reeb orbits of $\alpha_{0}$ are also $\mathbb{S}^{1}$-orbits, and have period $1$. In other words, $\alpha_{0}$ is a Zoll contact form on $M$, and as such has volume $|e|$ (see Proposition 3.3 in [30]). The volume of $\alpha$ can be written
$\text{Vol}(\alpha)=\int_{M}\alpha\wedge\mathrm{d}\alpha=\int_{M}(\tilde{K}^{\alpha})^{2}\alpha_{0}\wedge\mathrm{d}\alpha_{0}.$
Since $\tilde{K}^{\alpha}$ does not vanish, $(\tilde{K}^{\alpha})^{2}$ has a positive minimum that we write as $K_{0}^{2}$. This critical point of $\tilde{K}^{\alpha}$ corresponds to a closed Reeb orbit of period $|K_{0}|$, hence the systole of $\alpha$ is not larger than $|K_{0}|$. Finally, one has
$\text{Vol}(\alpha)=\int_{M}(\tilde{K}^{\alpha})^{2}\alpha_{0}\wedge\mathrm{d}\alpha_{0}\geq K_{0}^{2}\int_{M}\alpha_{0}\wedge\mathrm{d}\alpha_{0}=|e|K_{0}^{2}\geq|e|\text{sys}(\alpha)^{2}.$
Equality holds if and only $\tilde{K}^{\alpha}$ is constant, i.e. if and only if $\alpha$ is Zoll. ∎
Remark 2.5.
From now on, we will assume that $e>0$. The Giroux classification of tight invariant contact structure implies that for $\alpha$ in $\Omega(e)$, the set of Legendrian $\mathbb{S}^{1}$-orbits is non-empty and connected, in other words the regular level set $(K^{\alpha})^{-1}(0)$ is not empty and connected, and therefore a circle.
3 Invariant surfaces of section and potentials
This section contains the two main geometrical constructions underlying the proof of Theorem 1.2, namely an invariant surface of section, and a decomposition for invariant contact forms. These two constructions will allow us to introduce a family of potentials encoding most of the dynamic of the Reeb flow. We end this section by giving some important properties of these potentials.
3.1 An invariant surface of section
Let us fix $\alpha$ in $\Omega(e)$. In this section, we will construct an $\mathbb{S}^{1}$-invariant surface of section for the Reeb flow of $\alpha$. We write $Z$ for the regular level set $\tilde{K}^{-1}(0)\subset M$. Let $\mathcal{C}\subset\mathbb{S}^{2}$ be the union of the connected components of levels sets of $K$ that contain a critical point. It follows from the fact that $\mathrm{d}K$ doesn’t vanish on $\mathbb{S}^{2}\setminus\mathcal{C}$, that the difference $\mathbb{S}^{2}\setminus\mathcal{C}$ is a union of cylinders $(\Gamma_{i})_{i\in I}$. By Remark 2.5, $\pi(Z)$ is connected. Hence, there is a unique index $i_{0}$ such that $K$ has zeroes on $\Gamma_{i_{0}}$.
For each $i\in I$, we write $\epsilon_{i}^{+}$ and $\epsilon_{i}^{-}$ for the two components of $\partial\Gamma_{i}$, with $K(\epsilon_{i}^{+})>K(\epsilon_{i}^{-})$, and $K_{i}^{+}$ (resp. $K_{i}^{-}$) for $K(\epsilon_{i}^{+})$ (resp. $K(\epsilon_{i}^{-})$). We choose $x_{i}^{+}$ (resp. $x_{i}^{-}$) to be an arbitrary critical point of $K$ in $\epsilon_{i}^{+}$ (resp. $\epsilon_{i}^{-}$). Finally, we choose a curve $\gamma_{i}\colon]0,1[\rightarrow\Gamma_{i}$ such that:
  • $\gamma_{i}$ is transverse to the level sets of $K$;
  • $\lim_{s\rightarrow 0}\gamma_{i}(s)=x_{i}^{-}$ and $\lim_{s\rightarrow 1}\gamma_{i}(s)=x_{i}^{+}$.
Writing $\tilde{\gamma}_{i}$ for a lift of $\gamma_{i}$ to $M$, we set $\Sigma_{i}$ as the image of $\tilde{\gamma}_{i}$ under the $\mathbb{S}^{1}$ action, i.e. $\Sigma_{i}=\mathbb{S}^{1}\cdot\tilde{\gamma}_{i}=\pi^{-1}(\gamma_{i})$. Their union $\Sigma=\bigcup\limits_{i\in I}\Sigma_{i}$ is a disjoint union of cylinders. Its boundary components have the form $\mathbb{S}^{1}\cdot x$, with $x$ a critical point of $K$, and as such are both a closed Reeb orbit and an $\mathbb{S}^{1}$-orbit. The following proposition states that $\Sigma$ is a surface of section.
Proposition 3.1.
The vector field $R$ is transverse to $\Sigma$. Moreover, for all $x$ in $\Sigma$, the Reeb orbit of $x$ intersects $\Sigma$ in positive time.
Proof.
Let $x\in\Sigma$. Since the vector field $R$ is nowhere parallel to $\Theta$ on $M\setminus\pi^{-1}(\mathcal{C})$, its pushforward $\pi_{*}R$ is a nowhere vanishing vector field on $\mathbb{S}^{2}\setminus\mathcal{C}$. Moreover $\pi_{*}R$ is tangent to level sets of $K$, $K$ being perserved by the Reeb flow. Since $\gamma_{i}$ is transverse to the level sets of $K$, it is transverse to $\pi_{*}(R)$. Thus $R$ is transverse to $\Sigma=\pi^{-1}(\gamma_{i})$. The level sets of $K$ in $\mathbb{S}^{2}\setminus\mathcal{C}$ are unions of circles, hence the flow lines of $\pi_{*}R$ are circles as well. In particular, they pass infinitely many times through $\pi(x)$, which means that the Reeb orbit of $x$ in $M$ must intersect the $\mathbb{S}^{1}$-orbit of $x$, which is contained in $\Sigma_{i}$, infinitely many times. This proves the result. ∎
For $x$ in $\Sigma$, we will write $\tau(x)$ for the first positive time at which the Reeb orbit starting at $x$ hits $\Sigma$ again. This allows us to define a diffeomorphism $\Phi$ of $\Sigma$ as
$\begin{array}[]{ccccc}\Phi&\colon&\Sigma&\rightarrow&\Sigma\\ &&x&\mapsto&\phi^{\tau(x)}(x)\end{array}$
as well as $\Phi_{i}=\Phi_{|\Sigma_{i}}$. Note that periodic points of $\Phi$ correspond to closed Reeb orbits of $\alpha$.
3.2 A decomposition for invariant contact forms
In this section, we give a decomposition for invariant contact forms on $M$. We start with a preliminary lemma.
Lemma 3.2.
Let $U$ be an $\mathbb{S}^{1}$-invariant open set of $M$, and $\kappa$ a one-form on $U$. Then there is a one-form $\beta$ on $\pi(U)\subset\mathbb{S}^{2}$ such that $\kappa=\pi^{*}\beta$ if and only if $\mathcal{L}_{\Theta}\kappa=0$ and $i_{\Theta}\kappa=0$.
Proof.
The direct implication is easy. Conversely, let $\kappa$ satisfying $\mathcal{L}_{\Theta}\kappa=0$ and $i_{\Theta}\kappa=0$. We define $\beta$ by $\beta_{x}(X)=\kappa_{y}(Y)$, where $\pi(y)=x$ and $\mathrm{d}\pi_{y}(Y)=X$. Since $i_{\Theta}\kappa=0$, $\beta_{y}(Y)$ does not depend on the choice of $Y$. $\beta_{y}(Y)$ also does not depend on the choice of $y$ as $\mathcal{L}_{\Theta}\kappa=0$. ∎
Since $\mathbb{S}^{2}\setminus\pi(Z)$ is the union of two disks, and because the only $\mathbb{S}^{1}$-principal bundle over a disk is the trivial one, each of the two components of $M\setminus Z$ is a solid torus, the fundamental group of which being generated by the $\mathbb{S}^{1}$-orbits of the action. Let $\sigma$ be an $\mathbb{S}^{1}$-invariant closed one-form on $M\setminus Z$ such that $\int_{\mathbb{S}^{1}\cdot x}\sigma=1$ for all $x$ in $M\setminus Z$, the $\mathbb{S}^{1}\cdot x$ being oriented by the action. We will occasionally write $\sigma^{+}$ (resp. $\sigma^{-}$) for the restriction of $\sigma$ to $(K>0)$ (resp. to ($K<0$)).
Since the one-form $\alpha-K\sigma$ defined on $M\setminus Z$ is $\mathbb{S}^{1}$-invariant, and satisfies $i_{\Theta}(\alpha-K\sigma)=K-K=0$, according to Lemma 3.2 we have on $M\setminus Z$,
$\alpha=K\sigma+\pi^{*}\beta,$
where $\beta$ is a one-form on $\mathbb{S}^{2}\setminus\pi(Z)$.
We end this section with a useful geometrical interpretation of the Euler number $e$ of the $\mathbb{S}^{1}$-bundle $M\rightarrow\mathbb{S}^{2}$ (see for example [26] for a detailed discussion in the more general context of Seifert manifolds).
Remark 3.3.
[…]
We will write $c_{x}$ for the portion of Reeb orbit $\phi^{[0,\tau(x)]}$ starting at $x$ and ending at $\Phi(x)$. If $k\in]K^{-}_{i},K^{+}_{i}[$, there is a unique $x_{i,k}$ in $\tilde{\gamma}_{i}$ such that $K(x_{i,k})=k$. We will occasionally write $c_{i,k}$ for $c_{x_{i,k}}$.
Proposition 3.4.
The following identity holds:
$\int_{c_{i_{0},0}}\sigma^{+}-\sigma^{-}=e.$
Proof.
We write $D_{1}$ (resp. $D_{2}$) for the 2-disk $(K>0)$ (resp. $(K<0)$), and $\text{Crit}(f)$ for the set of critical values of a real-valued function $f$. The bundle $M$ is obtained by gluing the two bundles $T_{1}=\pi^{-1}(D_{1})$ and $T_{2}=\pi^{-1}(D_{2})$ via the attaching map described above. To understand the attaching map, we must understand the orientations of $\partial T_{1}$ and $\partial T_{2}$. Let $Y$ be an $\mathbb{S}^{1}$-invariant gradient-like vector field for $\tilde{K}$ on $M$, i.e. a vector field vanishing exactly at critical points of $\tilde{K}$, and such that $\mathrm{d}\tilde{K}(Y)>0$ otherwise. Since $\alpha\wedge\mathrm{d}\alpha(R,Y,X)=\mathrm{d}\alpha(Y,X)=\mathrm{d}K(Y)>0$ on $M\setminus\text{Crit}(\tilde{K})$, $(R,Y,X)$ is a positive basis on $M\setminus\text{Crit}(\tilde{K})$. Thus, $(\pi_{*}R,\pi_{*}Y)$ is a positive basis on $\mathbb{S}^{2}\setminus\text{Crit}(K)$. We get that the induced boundary orientation of $\partial D_{1}$ (resp. of $\partial D_{2}$) coincides with (resp. is the opposite of) the orientation induced by the Reeb flow. Following Remark 3.3, the attaching map $\phi$ acts on the level of fundamental groups via the map $\phi_{*}=\begin{pmatrix}1&-e\\ 0&-1\end{pmatrix}$ in the basis $(h_{1},\mu_{1})$ and $(h_{2},\mu_{2})$. Hence, the induced map $\phi^{*}\colon H^{1}(\partial T_{2},\mathbb{Z})\rightarrow H^{1}(\partial T_{1},\mathbb{Z})$ is $\begin{pmatrix}1&0\\ -e&-1\end{pmatrix}$ in the dual basis $(h_{1}^{*},\mu_{1}^{*})$ and $(h_{2}^{*},\mu_{2}^{*})$. Since $\sigma^{+}=h_{1}^{*}$ on $T_{1}$ and $\sigma^{-}=h_{2}^{*}$ on $T_{2}$, and since $\pi(c_{i_{0},0})=\mu_{1}$ in $\pi_{1}(\partial T_{1})$, we get
$\int_{c_{i_{0},0}}\sigma^{+}-\sigma^{-}=\int_{c_{i_{0},0}}h_{1}-\phi^{*}{h_{2}}=\int_{\mu_{1}}e\mu_{1}^{*}=e.$
3.3 A family of potentials
For each $i\in I$ we define the map
$\begin{array}[]{ccccc}J_{i}&\colon&]K_{i}^{-},K_{i}^{+}[&\rightarrow&\mathbb{R}\\ &&k&\mapsto&\int_{\pi(c_{i,k})}\beta.\end{array}.$
Since the curves $\pi(c_{i,k})$ are closed curves that uniformly converge to $\epsilon_{i}^{\pm}$ when $k$ approaches to $K_{i}^{\pm}$, $J_{i}$ can be extended continuously to $[K^{-}_{i},K^{+}_{i}]$. Moreover, since $\lim\limits_{k\rightarrow 0,k>0}J_{i_{0}}(k)=\lim\limits_{k\rightarrow 0,k>0}\int_{c_{i,k}}\alpha-k\sigma=\int_{c_{i_{0},0}}\alpha$, and similarly $\lim\limits_{k\rightarrow 0,k<0}J_{i_{0}}(k)=\int_{c_{i_{0},0}}\alpha$, $J_{i_{0}}$ is continuous at 00.
Lemma 3.5.
The map $J_{i}$ is differentiable on $]K_{i}^{-},K_{i}^{+}[\setminus\{0\}$, with $J_{i}^{\prime}(k)=-\int_{c_{i,k}}\sigma$. In addition,
$J^{\prime}_{i_{0}}(0^{+})-J^{\prime}_{i_{0}}(0^{-})=-e.$
Proof.
Let $k^{-}<k^{+}$ be in $]K_{i}^{-},K_{i}^{+}[$, such that $k^{-}$ and $k^{+}$ have same sign. Define $S$ to be $\bigcup\limits_{k^{-}<k<k^{+}}c_{i,k}$. This surface projects injectively on $\mathbb{S}^{2}$, with image $\Gamma_{i}\cap K^{-1}(]k^{-},k^{+}[)$, whose oriented boundary is $\pi(k^{-})-\pi(k^{+})$ (both oriented by the flow of $\pi_{*}R$). By Stokes theorem, one then has
$J_{i}(k^{+})-J_{i}(k^{-})$ $=-\int_{\pi(S)}\mathrm{d}\beta$
$=-\int_{S}\pi^{*}\mathrm{d}\beta$
$=-\int_{S}\mathrm{d}\alpha-\mathrm{d}K\wedge\sigma$
$=-\int_{S}\sigma\wedge\mathrm{d}K$
$=\int_{k^{-}}^{k^{+}}\left(-\int_{c_{i,k}}\sigma\right)\mathrm{d}k.$
The penultimate equality follows from the fact that the restriction of $\mathrm{d}\alpha$ to $TS$ vanishes as $R$ is tangent to $S$, and the last equality is obtained after the change of variable $x\in\tilde{\gamma}_{i}\cap S\mapsto K(x)\in]k^{-},k^{+}[$. Dividing by $k^{+}-k^{-}$ and letting $k^{+}$ go to $k^{-}$ yields the result. Finally, $J^{\prime}_{i_{0}}(0^{+})-J^{\prime}_{i_{0}}(0^{-})=-e$ follows from Proposition 3.4. ∎
An immediate corollary of Lemma 3.5 is the below identity relating $\tau$ and $J_{i}$.
Corollary 3.6.
For all $k\in]K_{i}^{-},K_{i}^{+}[$, $\tau(x_{i,k})=J_{i}(k)-kJ_{i}^{\prime}(k)$.
Proof.
Integrating the identity $\alpha=K\sigma+\pi^{*}\beta$ along $c_{i,k}$ yields
$\tau(x_{i,k})=k\int_{c_{i,k}}\sigma+J_{i}(k).$
By Lemma 3.5, $\int_{c_{i,k}}\sigma=-J_{i}^{\prime}(k)$ and the result follows. ∎
A second corollary of Lemma 3.5 is that the derivative of $J_{i}$ is closely related to periodic points of the first-return map of the surface of section.
Corollary 3.7.
The Reeb orbit starting at $x_{i,k}\in\tilde{\gamma}_{i}$ is closed if and only if $J_{i}^{\prime}(k)$ is rational. Moreover, if $J_{i}^{\prime}(k)=\frac{p}{q}$ with $\gcd(p,q)=1$, then $x_{i,k}$ is a $q$-periodic point of $\Phi_{i}$, and the corresponding closed Reeb orbit has minimal period $q(J_{i}(k)-kJ_{i}^{\prime}(k))$.
Proof.
Since $\tilde{K}$ is preserved by the Reeb flow, the first-return map $\Phi_{i}$ of $\Sigma_{i}$ induces a circle diffeomorphism of $\mathbb{S}^{1}\cdot x_{i,k}$ for each $k\in]K_{i}^{-},K_{i}^{+}[$. By $\mathbb{S}^{1}$-invariance of the Reeb flow, each of those circle diffeomorphism is a translation. Finally, for $k\neq 0$, homotoping $c_{i,k}$ relative to its boundary in $M\setminus Z$ to a path contained in $\mathbb{S}^{1}\cdot x_{i,k}$, together with the closedness of $\sigma$, gives us that the circle diffeomorphism of $\mathbb{S}^{1}\cdot x_{i,k}$ is a translation of shift $\int_{c_{i,k}}\sigma$. As a consequence, the Reeb orbit starting at $x_{i,k}$ is closed if and only if $\int_{c_{i,k}}\sigma=-J_{i}(k)$ is a rational number.
When $J_{i}^{\prime}(k)=\frac{p}{q}$ with $\gcd(p,q)=1$, the translation of shift $\int_{c_{i,k}}\sigma=-\frac{p}{q}$ is $q$-periodic. Furthermore, the closed Reeb orbit starting at $x_{i,k}$ has period $q\tau(x_{i,k})=q(J_{i}(k)-kJ_{i}^{\prime}(k))$ according to Corollary 3.6. ∎
We now give some properties of the family of potentials $(J_{i})_{i\in I}$. First of all, Lemma 3.8 tells us that the integrals of the $(J_{i})_{i\in I}$ are related to the volume.
Lemma 3.8.
The identity
[…]
holds.
Proof.
Since $\sigma$ is closed, the identity $\alpha=K\sigma+\pi^{*}\beta$ implies that
$\alpha\wedge\mathrm{d}\alpha=\sigma\wedge\pi^{*}(K\mathrm{d}\beta+\beta\wedge\mathrm{d}K).$
Let $T$ be a connected component of $M\setminus Z$. The volume of $T$ can be written as
$\text{Vol}(T)$ $=\int_{T}\sigma\wedge\pi^{*}(K\mathrm{d}\beta+\beta\wedge\mathrm{d}K)$
$=\int_{\pi(T)}K\mathrm{d}\beta+\beta\wedge\mathrm{d}K$
$=\int_{\pi(T)}\mathrm{d}(K\beta)+2\beta\wedge\mathrm{d}K$
$=\int_{\pi(Z)}K\beta+2\int_{\pi(T)}\beta\wedge\mathrm{d}K$
$=2\int_{\pi(T)}\beta\wedge\mathrm{d}K$
$=2\int_{\pi(T)\setminus\mathcal{C}}\beta\wedge\mathrm{d}K$
$=2\sum\limits_{i\in I}\int_{\Gamma_{i}\cap\pi(T)}\beta\wedge\mathrm{d}K$
$=2\sum\limits_{i\in I}\int_{]K_{i}^{-},K_{i}^{+}[\cap K(\pi(T))}\left(\int_{\pi(c_{i,k})}\beta\right)\mathrm{d}k.$
The last equality follows from the change of variable $x\in\gamma_{i}\cap\pi(T)\mapsto K(x)\in]K_{i}^{-},K_{i}^{+}[\cap K(\pi(T))$. The claimed identity is then obtained by summing over both components $T$ of $M\setminus Z$. ∎
The key properties of the family $(J_{i})_{i\in I}$ all follow from a differential inequality involving $J_{i}$.
Proposition 3.9.
For all $i$ in $I$, the map $J_{i}\colon]K_{i}^{-},K_{i}^{+}[\setminus\{0\}\rightarrow\mathbb{R}$ satisfies
$\forall k\in]K_{i}^{-},K_{i}^{+}[\setminus\{0\},J_{i}(k)-kJ_{i}^{\prime}(k)>0.$
Proof.
This follows immediately from Corollary 3.6 in addition to the positivity of $\tau$. ∎
We now state some properties of this differential inequality.
Lemma 3.10.
[…]
$f(x)-xf^{\prime}(x)>0.$ (1)
[…]
Proof.
Without loss of generality, we can assume that $f(b)\geq 0$ and $0<a<b$. If there is an $x_{0}\in]a,b[$ such that $f(x_{0})\leq 0$, it follows immediately from the inequality (1), plus the fact that $x_{0}>0$, that $f^{\prime}(x_{0})<0$. Furthermore, the inequality (1) implies that $f(x)>0$ whenever $f^{\prime}(x)=0$. Thus we get that $f^{\prime}$ is negative on $[x_{0},b[$. As a consequence, $f(x)<0$ on $[x_{0},b]$ and in particular $f(b)<0$, which is a contradiction. ∎
Lemma 3.11.
The function
[…]
is positive.
Proof.
Let $T$ be a component of $M\setminus Z$. Assume without loss of generality that $\tilde{K}$ is positive on $T$. We start by proving the result when $K$ has isolated critical points.
  • When $K$ has isolated critical points: Let $I_{T}=\{i\in I,\tilde{\gamma}_{i}\cap T\neq\emptyset\}$, and $i\in I_{T}$. We claim that when $K$ has isolated critical points, $f$ extends continuously to $[\min(K),\max(K)]$. Indeed, in that case if $k_{0}\in\text{Crit}(K)$, the curve $K^{-1}(k)$ converges to $K^{-1}(k_{0})$ for the Hausdorff topology when $k$ goes to $k_{0}$. As a consequence, $f(k)$ converges to $\int_{K^{-1}(k_{0})}\beta$ when $k$ goes to $k_{0}$. Let us write $K_{1}<\ldots<K_{p}=\max K$ for the ordered critical values of $K$ on $T$, and set $K_{0}=0$. $f$ is smooth on $]0,K_{p}]\setminus\{K_{1},\ldots,K_{p}\}$, and continuous on $[0,K_{p}]$. Moreover, $f_{|]K_{l},K_{l+1}[}$ is the sum of all the $J_{i}$ such that $]K_{l},K_{l+1}[\subset]K_{i}^{-},K_{i}^{+}[$. By Proposition 3.9 applied to each of those $J_{i}$ and by linearity, $f$ satisfies the differential inequality
    $f(x)-xf^{\prime}(x)>0$
    on each $]K_{l},K_{l+1}[$. Since $K^{-1}(K_{p})$ consists of isolated points, all of them maximizers for $K$, $f(K_{p})$ is zero. Lemma 3.10 gives that $f$ is positive on $]K_{p-1},K_{p}[$. Moreover, continuity of $f$ at every $K_{i}$ when $i<p$ allows us to apply Lemma 3.10 on each $]K_{p-l-1},K_{p-l}[$ successively when $l$ goes from $1$ to $p-1$. This gives us the result.
  • The general case: Let $k\in[\min(K),\max(K)]\setminus\text{Crit}(K)$. Since $[\min(K),\max(K)]\setminus\text{Crit}(K)$ is open, let $[k-\eta,k+\eta]$ be a closed neighborhood of $k$ in $[\min(K),\max(K)]\setminus\text{Crit}(K)$, and let $\tilde{g}\colon M\rightarrow]0,+\infty[$ a smooth function such that:
    • $\tilde{g}$ is $\mathbb{S}^{1}$-invariant and hence induces a function $g\colon\mathbb{S}^{2}\rightarrow]0,+\infty[$;
    • $\tilde{g}$ is constant equal to one on $\tilde{K}^{-1}([k-\eta,k+\eta])$;
    • $gK$ has isolated critical points.
    Such a function $\tilde{g}$ exists since $[k-\eta,k+\eta]$ only contains regular values of $K$. Then $\tilde{g}\alpha$ is an $\mathbb{S}^{1}$-invariant contact form on $M$, with $\tilde{K}^{\tilde{g}\alpha}=\tilde{g}\tilde{K}$. Since $gK$ has isolated critical points, the previous argument implies that $f^{\tilde{g}\alpha}$ is positive on $[\min(gK),\max(gK)]\setminus\text{Crit}(gK)$. Since $\tilde{g}$ is constant equal to one on $\tilde{K}^{-1}([k-\eta,k+\eta])$, $K^{\tilde{g}\alpha}=K$ and $\beta^{\tilde{g}\alpha}=\beta^{\alpha}$ on $K^{-1}([k-\eta,k+\eta])$, we have $f^{\tilde{g}\alpha}=f^{\alpha}$ on $[k-\eta,k+\eta]$. In particular, $f^{\alpha}(k)>0$. The result then follows as $k$ is an arbitrary regular value of $K$.
4 Proof of Theorem 1.2
In this section, we prove Theorem 1.2. Since the case of $\mathbb{S}^{1}$-actions transverse to the contact structure has been taken care of in Proposition 2.4, we focus now on the case where $e>0$ and for which the set of set of Legendrian $\mathbb{S}^{1}$-orbits is non-empty and connected .
Proof.
(Theorem 1.2) We claim it is enough to prove the result for $e=1$. Indeed, $M$ has universal cover $\mathbb{S}^{3}$, and the covering map $\mathbb{S}^{3}\rightarrow M$ has degree $e$. Lifting the $\mathbb{S}^{1}$-action and $\alpha$ by this covering map yields a free action of $\mathbb{S}^{1}$ on $\mathbb{S}^{3}$ of Euler number 1 and an invariant contact form $\tilde{\alpha}\in\Omega(1)$. This contact form has volume $\text{Vol}(\tilde{\alpha})$ equal to $e\text{Vol}(\alpha)$. Moreover, a closed Reeb orbit of $\alpha$ with period $T$ lifts to a closed Reeb orbit of $\tilde{\alpha}$ with period $eT$. In particular, we have $\text{sys}(\tilde{\alpha})=e\hskip 2.84544pt\text{sys}(\alpha)$. Finally, $\alpha$ is Zoll if and only if $\tilde{\alpha}$ is Zoll. The claim follows.
Let $\alpha$ be in $\Omega(1)$. We write $J$ for $J_{i_{0}}$, and $K_{\text{min}}$ for the minimal absolute value of a critical value of $K$. Since $]-K_{\text{min}},K_{\text{min}}[$ is contained in $U_{i_{0}}$ and does not intersect $]K_{i}^{-},K_{i}^{+}[$ when $i\neq i_{0}$, Lemma 3.11 tells us that $J$ is positive on $[-K_{\text{min}},K_{\text{min}}]$. By Lemma 3.8 and 3.11, we have
$\text{Vol}(\alpha)\geq 2\int_{-K_{\text{min}}}^{K_{\text{min}}}J\mathrm{d}k.$ (2)
Additionally, the following holds:
$\begin{split}\int_{-K_{\text{min}}}^{K_{\text{min}}}J\mathrm{d}k&=\int_{-K_{\text{min}}}^{K_{\text{min}}}\left(J+\frac{|k|}{2}\right)\mathrm{d}k-\int_{-K_{\text{min}}}^{K_{\text{min}}}\frac{|k|}{2}\mathrm{d}k\\ &=\int_{-K_{\text{min}}}^{K_{\text{min}}}\left(J+\frac{|k|}{2}\right)\mathrm{d}k-\frac{K_{\text{min}}^{2}}{2}.\end{split}$ (3)
We write $g$ for the function $k\in[-K_{\text{min}},K_{\text{min}}]\mapsto J(k)+\frac{|k|}{2}$. It is continuously differentiable on $]-K_{\text{min}},K_{\text{min}}[$ by virtue of Lemma 3.5 and the assumption $e=1$. We claim that the lower bound
$g(k)\geq\max\left(\frac{|k|}{2},\frac{\text{sys}(\alpha)}{2}\right)$ (4)
holds. Indeed, since $J$ is positive, one has that $g(k)\geq\frac{|k|}{2}$. Furthermore, we have
$g(\pm K_{\text{min}})\geq\frac{K_{\text{min}}}{2}\geq\frac{\text{sys}(\alpha)}{2}.$ (5)
Finally, if $g$ reaches its minimum at a critical point $k_{0}\in]-K_{\text{min}},K_{\text{min}}[$, then:
  • if $k_{0}\neq 0$, we have $|J^{\prime}(k_{0})|=\frac{1}{2}$;
  • if $k_{0}=0$, then for $g$ to be minimal at 0, we must have that $J(k)+\frac{1}{2}k\geq J(0)$ for $k$ positive, implying $J^{\prime}(0^{+})\geq-\frac{1}{2}$. Similarly, we must have $J(k)-\frac{1}{2}k\geq J(0)$ for $k$ negative, which gives us $J^{\prime}(0^{-})\leq\frac{1}{2}$. Since $J^{\prime}(0^{+})-J^{\prime}(0^{-})=-1$, we get that $J^{\prime}(0^{+})=-\frac{1}{2}$ and $J^{\prime}(0^{-})=\frac{1}{2}$.
By Corollary 3.7, $k_{0}$ corresponds to a $2$-periodic point of the first-return map of the surface of section, hence a periodic orbit of the Reeb flow, of period twice the corresponding critical value of $g$. We get that critical values of $g$ are not smaller than $\frac{\text{sys}(\alpha)}{2}$. So, the minimum of $g$ being either reached at ±$\pm K_{\text{min}}$ or at a critical point, we get by (5) that
$\min(g)\geq\frac{\text{sys}(\alpha)}{2}.$
The claim (4) follows. Integrating the lower bound (4) in (3) and substituting it in (2) gives that
$\text{Vol}(\alpha)$ $\geq 2\int_{-K_{\text{min}}}^{K_{\text{min}}}J\mathrm{d}k$
$\geq\int_{-K_{\text{min}}}^{K_{\text{min}}}\max\left(|k|,\text{sys}(\alpha)\right)\mathrm{d}k-K_{\text{min}}^{2}$
$\geq\text{sys}(\alpha)^{2}+K_{\text{min}}^{2}-K_{\text{min}}^{2}$
$\geq\text{sys}(\alpha)^{2}.$
This returns the wanted inequality for $e=1$. Moreover, the equality case is obtained if and only if equality holds in (2) and (4). Equality holds in (3) if and only if $K$ has exactly two critical values $K_{\text{min}}$ and $-K_{\text{min}}$. In particular, all Reeb orbits that do not intersect the surface of section $\Sigma$ are closed, with period $K_{\text{min}}$. Furthermore, equality holds in (4) if and only if $\text{sys}(\alpha)$ equals $K_{\text{min}}$, since otherwise $g$ would not be smooth on $]-K_{\text{min}},K_{\text{min}}[\setminus\{0\}$. This happens if and only if $g$ is constant equal to $\frac{K_{\text{min}}}{2}$, i.e. $J(k)=\frac{K_{\text{min}}}{2}-\frac{|k|}{2}$. In that case, $|J^{\prime}(k)|=\frac{1}{2}$ for all $k$, thus $x_{i_{0},k}$ is a $2$-periodic point of the surface of section. The Reeb orbit starting at any of the $x_{i_{0},k}$ is then closed, with minimal period $2\tau(k)=2(J(k)-kJ^{\prime}(k))=K_{\text{min}}$. We getthat every Reeb orbit of $\alpha$ is closed, with common minimal period $K_{\text{min}}$. In other words, $\alpha$ is Zoll. ∎